# The category of anyon sectors for non-abelian quantum double models

### v1

@@ 1 Introduction
Kitaev’s quantum double model [kitaev2003fault] is the prototypical example of a topologically ordered quantum spin system with long-range entanglement (see [MR3929747] for an introduction). Such models host quasi-particle excitations with non-trivial braid statistics called anyons. The physical properties of such anyons (such as their behavior under exchange or fusion) can be described algebraically by braided (and often even modular) tensor categories [Kitaev2006, Wang2010]. In this paper we show that for the quantum double model for a finite gauge group $G$, defined on the plane, this briaded tensor category can be recovered from the unique frustration-free ground state of the model (under some mild technical assumption), and is given by $\mathbf{Rep}_{f}{\mathcal{D}}(G)$, the category of finite dimensional unitary representations of the quantum double algebra of $G$.
Our approach is motivated by the Doplicher–Haag–Roberts (DHR) theory of superselection sectors (see [HaagLQP] for an overview). Mathematically, we can identify the anyons with certain equivalence classes of irreducible representations of the (quasi-local) observable algebra $\mathcal{A}$.
@@
This intuition is conveniently captured by the superselection criterion. Namely, a representation $\pi$ satisfies the superseleciton criterion if
@@
where $\Lambda$ is any cone (a notion which we will make more precise later) and $\Lambda^{c}$ is its complement, $\pi_{0}$ is the GNS representation of the (unique) frustration free ground state of the quantum double model, and $\mathcal{A}_{\Lambda^{c}}$ is the ${\rm C}^{*}$-algebra generated by all local observables localized in $\Lambda^{c}$. That is, we consider representations that, outside any cone, are unitarily equivalent to the ground state representation. A superselection sector (or simply anyon sector) is an equivalence class of such representations.
They key insight of Doplicher, Haag and Roberts is that the superselection sectors are naturally endowed with a monoidal product (‘fusion’) and a symmetry describing the exchange of bosonic/fermionic sectors. This was later extended to describe braiding statistics [frs1, frs2], yielding a braided monoidal category. These categories precisely capture the physical properties of anyon sectors, including their braiding and fusion rules. The key technical step is that, using a technical property called Haag duality, one can pass from representations to endomorphisms of the quasi-local algebra which are localized (i.e., they act non-trivially only in the localization region) and transportable (the localization region can be moved around with unitaries). See [HalvorsonMueger] for an overview of this construction in the language of $\rm{C}^{*}$-tensor categories. This theory was initially developed in the context of relativistic quantum field theories. The construction has later been adapted to quantum spin systems, see e.g. [Naaijkens2011, Fiedler2014, Ogata2021]. For a recent completely axiomatic approach towards anyon sector theory, see [Bhardwaj2024].
In this paper we study the anyon sector theory, including fusion and braiding rules, of the quantum double model for arbitrary finite gauge group $G$ [kitaev2003fault], extending previous results obtained for abelian $G$ [Fiedler2014]. In particular, our main result can be paraphrased as follows:
@@
We will give a precise statement of our main result (including our assumptions) later when we have introduced the necessary terminology, but remark that Haag duality for cones is a technical property that holds for the abelian quantum double model [Fiedler2014], and one can still construct a category of anyon sectors without it (or with a weaker version thereof). See Remark 2.3 below for more details. We also note that since $\mathbf{Rep}_{f}{\mathcal{D}}(G)$ is a unitary modular tensor category, the category of anyon sectors is as well.
@@
The main idea behind our proof is as follows. For each irreducible represention of ${\mathcal{D}}(G)$, examples of representations $\pi$ satisfying the superselection criterion (1.1) were constructed in [Naaijkens2015]. It was then shown in [bols2023classificationanyonsectorskitaevs] that these representations are irreducible, and in fact form a complete set of representatives of irreducible representations satisfying (1.1). These irreducible anyon sectors correspond to the simple objects (i.e., the anyon types) in our category. Our main contribution in this paper is to establish the full superselection structure, including fusion and braiding.
@@
The key difference with the abelian case studied in [Naaijkens2011, Fiedler2014] is the use of amplimorphisms, i.e. *-homomorphisms $\chi:\mathcal{A}\to M_{n}(\mathcal{A})$, instead of endomorphisms. [^1] For technical reasons we will in fact need to consider amplimorphisms of some slightly bigger algebra $\mathcal{B}$ containing $\mathcal{A}$. This can be understood as follows: recall that in the quantum double models, we can define ‘ribbon operators’ which create a pair of excitations from the ground state. To obtain single-anyon states, one sends one of the excitations off to infinity. For each irrep of ${\mathcal{D}}(G)$, there is a corresponding multiplet of ribbon operators, transforming according to the irrep, with the total number of operators in the multiplet given by the dimension of the irrep. Hence for non-abelian representations, one has more than one ribbon operator, which combine naturally into an amplimorphism.
@@
Although it is possible to pass from amplimorphisms to endomorphisms, as we shall see later, doing so requires making some choices, and one loses the explicit description of the map. Hence to identify the full superselection theory, we work mainly in the amplimorphism picture. In particular, we show that the amplimorphisms constructed in [Naaijkens2015] can be endowed with a tensor product and a braiding, analogous to the tensor product and braiding of endomorphisms in the DHR theory. More precisely, we construct a braided ${\rm C^{*}}$-tensor category $\mathbf{Amp}$ of localized and transportable amplimorphisms, which includes as objects the amplimorphisms constructed in [Naaijkens2015]. We then consider the full subcategory $\mathbf{Amp}_{f}$ of $\mathbf{Amp}$ whose objects $\chi$ have finite dimensional $\mathrm{Hom}(\chi,\chi)$. This category can be shown to be semi-simple and closed under the monoidal product on $\mathbf{Amp}$, and we study the fusion rules (how tensor products decompose into irreducible objects) and the braiding. The result is that the category $\mathbf{Amp}_{f}$ of such amplimorphisms is equivalent to $\mathbf{Rep}_{f}{\mathcal{D}}(G)$ as braided tensor categories. Using the classification result of anyon sectors in this model obtained by two of the authors [bols2023classificationanyonsectorskitaevs], it then follows that the list of constructed anyon sectors is a complete list of representatives of irreducible anyon sectors. This then completes the classification.
@@
A similar approach using amplimorphisms was taken in [SzlachanyiV93, NillS97] to analyze topological defects of certain 1D quantum spin systems. In their setting the anyon sectors are localized in finite intervals, with the corresponding algebra of observables localized in that region being finite dimensional. This necessitated the use of amplimorphisms instead of endomorphisms. In our case localization is in infinite cone regions, and the situation is different. In particular, the unitary operators that can move the localization regions around no longer live in the quasi-local algebra $\mathcal{A}$. From a technical point of view this means that we cannot restrict to a purely $\rm{C}^{*}$-algebraic approach with operators in the quasi-local algebra (or suitable amplifications) only, but have to consider von Neumann algebras as well, in particular the cone algebras $\pi_{0}(\mathcal{A}_{\Lambda})^{\prime\prime}$. [^2] This is already true for the abelian case, it is not specific to the non-abelian model. These cone algebras are “big enough” in the sense that they are properly infinite [Naaijkens2012, Fiedler2014, Tomba2023]. This allows us to directly relate the localized and transportable amplimorphisms to localized and transportable endomorphisms of some suitably defined auxiliary algebra, making the connection with the usual DHR theory in terms of endomorphisms.
@@
The paper is outlined as follows. In Section 2 we define the quantum double model and the associated categories of localized and tranportable amplimorphisms $\mathbf{Amp}$ and endomorphisms $\mathbf{DHR}$, as well as their ‘finite’ versions $\mathbf{Amp}_{f}$ and $\mathbf{DHR}_{f}$. We then state our main theorem, namely that the categories $\mathbf{Amp}_{f}$ and $\mathbf{DHR}_{f}$ are braided $\rm C^{*}$-tensor categories, equivalent to $\mathbf{Rep}_{f}{\mathcal{D}}(G)$. Section 3 is devoted to spelling out the braided $\rm C^{*}$-tensor structure of $\mathbf{Amp}$ and $\mathbf{DHR}$. These two categories are then shown to be equivalent in Section 4. Explicit localized and tranportable amplimorphisms corresonding to representations of ${\mathcal{D}}(G)$ are constructed in Section 5 by taking limits of ‘ribbon multiplets’. These explicit amplimorphisms are organized into full subcategories $\mathbf{Amp}_{\rho}$ of $\mathbf{Amp}$ for a fixed half-infinite ribbon $\rho$, which are later shown to be equivalent to $\mathbf{Amp}_{f}$. This section also establishes the key properties of these ribbon multiplets that underlie the fusion and braiding structure of $\mathbf{Amp}_{f}$. In Section 6 we rephrase the main result of [bols2023classificationanyonsectorskitaevs], namely that the amplimorphisms corresponding to irreducible representations of ${\mathcal{D}}(G)$ constructed in the previous section exhaust all simple objects of $\mathbf{Amp}$. Together with semi-simplicity of $\mathbf{Amp}_{f}$, this implies that the $\mathbf{Amp}_{\rho}$ are full and faithful subcategories of $\mathbf{Amp}_{f}$. Finally, Section 7 proves the main theorem. The appendices collect well-known facts about ribbon operators and some technical results related to taking their limits.
@@ 2.1 The quantum double model and its ground state
We first recall the definition of the quantum double model [kitaev2003fault] and introduce our notation. Throughout the paper, we fix a finite group $G$. Let $\Gamma$ be the triangular lattice in ${\mathbb{R}}^{2}$ and denote by $\Gamma^{E}$ the collection of oriented edges of $\Gamma$ which are oriented towards the right, see Figure 1.333We use the triangular lattice for simplicity, and to work in the same setting as [bols2023classificationanyonsectorskitaevs], but believe the results hold for more general planar graphs as well. Denote by $\Gamma^{V},\Gamma^{F}$ the set of vertices and faces of $\Gamma$ respectively. To each edge $e\in\Gamma^{E}$ we associate a degree of freedom ${\mathcal{H}}_{e}\simeq{\mathbb{C}}[G]$ with basis $\{|g\rangle_{e}\,:\,g\in G\}$ and corresponding algebra ${\mathcal{A}}_{e}=\mathrm{End}({\mathcal{H}}_{e})\cong M_{|G|}(\mathbb{C})$. We define in the usual way local algebras of observables ${\mathcal{A}}^{\mathrm{loc}}_{X}$ supported on any $X\subset\Gamma^{E}$ and their norm closures ${\mathcal{A}}_{X}:=\overline{{\mathcal{A}}_{X}^{\mathrm{loc}}}^{\|\cdot\|}$. We write ${\mathcal{A}}={\mathcal{A}}_{\Gamma^{E}}$ and ${\mathcal{A}}^{\mathrm{loc}}={\mathcal{A}}_{\Gamma^{E}}^{\mathrm{loc}}$.
@@
The following theorem is proven in various sources [Fiedler2014, Cui2020kitaevsquantum, Tomba2023, bols2023classificationanyonsectorskitaevs].
@@ 2.2 Cone algebras, Haag duality, and the allowed algebra
We remark that all these cone algebras are infinite factors [Naaijkens2012, Ogata2021]. We will moreover assume that Haag duality holds for cones.
@@
Haag duality for cones is proven in [Fiedler2014] in the case $G$ is an abelian group. We expect it to hold in the non-abelian case as well. It is possible to only demand the weaker condition of approximate Haag duality (see [Ogata2021]), but this makes the analysis more technical and hence we prefer to assume the above form.
@@
is the $\rm C^{*}$-algebra generated by the cone algebras of allowed cones. Note that the set of allowed cones is a directed set for the inclusion relation.
@@ 2.3 Categories of amplimorphisms and endomorphisms
A *-homomorphism $\chi:{\mathcal{B}}\rightarrow M_{n\times n}({\mathcal{B}})$ is called an amplimorphism of degree $n$. We do not require such amplimorphisms to be unital. Given two amplimorphisms $\chi$ and $\chi^{\prime}$ of degrees $n$ and $n^{\prime}$ respectively, we let
@@
In the algebraic description of anyons, it is commonly assumed that all anyons have a conjugate (see for example  [Wang2010, Sect. 6.3]), meaning that each anyon type can fuse to the vacuum with some conjugate type.
The assumption that an object in a $\rm C^{*}$-tensor category has a conjugate implies that it has a finite-dimensional endomorphism space [LongoRoberts97, Lemma 3.2]. This is another way to see the necessity of restricting our attention to $\mathbf{Amp}_{f}$ if we want to show equivalence with $\mathbf{Rep}_{f}{\mathcal{D}}(G)$. Indeed, all finite dimensional representations of ${\mathcal{D}}(G)$ have conjugates.
@@
$\mathbf{DHR}$ is a braided $\rm C^{*}$-tensor subcategory of $\mathbf{Amp}$, see Section 3. We show in Section 7 that $\mathbf{DHR}_{f}$ is closed under the monoidal product of $\mathbf{DHR}$ and therefore inherits the braided $\rm C^{*}$-tensor structure.

### v2

@@ 1 Introduction
Kitaev’s quantum double model [Kit03] is the prototypical example of a topologically ordered quantum spin system with long-range entanglement (see [Zen+19] for an introduction). Such models host quasi-particle excitations with non-trivial braid statistics called anyons. The physical properties of such anyons (such as their behavior under exchange or fusion) can be described algebraically by braided (and often even modular) tensor categories [Kit06, Wan10]. In this paper we show that for the quantum double model for a finite gauge group $G$, defined on the plane, this braided tensor category can be recovered from the unique frustration-free ground state of the model (under some mild technical assumption), and is given by $\mathbf{Rep}_{f}{\mathcal{D}}(G)$, the category of finite dimensional unitary representations of the quantum double algebra of $G$.
Our approach is motivated by the Doplicher–Haag–Roberts (DHR) theory of superselection sectors (see [Haa96] for an overview). Mathematically, we can identify the anyons with certain equivalence classes of irreducible representations of the (quasi-local) observable algebra $\mathcal{A}$.
@@
This intuition is conveniently captured by the superselection criterion. Namely, a representation $\pi$ satisfies the superselection criterion if
@@
[…]
The key insight of Doplicher, Haag and Roberts is that the superselection sectors are naturally endowed with a monoidal product (‘fusion’) and a symmetry describing the exchange of bosonic/fermionic sectors. This was later extended to describe braiding statistics [FRS89, FRS92], yielding a braided monoidal category. These categories precisely capture the physical properties of anyon sectors, including their braiding and fusion rules. The essential technical step is that, using a technical property called Haag duality, one can pass from representations to endomorphisms of the quasi-local algebra which are localized (i.e., they act non-trivially only in the localization region) and transportable (the localization region can be moved around with unitaries). See [HM06] for an overview of this construction in the language of $\rm{C}^{*}$-tensor categories. This theory was initially developed in the context of relativistic quantum field theories. The construction has later been adapted to quantum spin systems, see e.g. [Naa11, FN15, Oga22]. For a recent completely axiomatic approach towards anyon sector theory, see [Bha+25].
In this paper we study the anyon sector theory, including fusion and braiding rules, of the quantum double model for arbitrary finite gauge group $G$ [Kit03], extending previous results obtained for abelian $G$ [FN15]. In particular, our main result can be paraphrased as follows:
@@
We will give a precise statement of our main result (including our assumptions) later when we have introduced the necessary terminology, but remark that Haag duality for cones is a technical property that holds for the abelian quantum double model [FN15], and one can still construct a category of anyon sectors without it (or with a weaker version thereof). A proof of Haag duality for a large class of models has recently been announced [OPR25]. See Remark 2.3 below for more details. We also note that since $\mathbf{Rep}_{f}{\mathcal{D}}(G)$ is a unitary modular tensor category, the category of anyon sectors is as well.
@@
As mentioned earlier, our assumptions imply that there is a braided $\rm{C}^{*}$-category of superselection sectors [FN15, Oga22]. Our main contribution in this paper is to construct this category explicitly for the quantum double model for all finite groups $G$. The main idea is as follows. For each irreducible represention of ${\mathcal{D}}(G)$, examples of representations $\pi$ satisfying the superselection criterion (1.1) were constructed in [Naa15]. It was then shown in [BV25] that these representations are irreducible, and in fact form a complete set of representatives of irreducible representations satisfying (1.1). These irreducible anyon sectors correspond to the simple objects (i.e., the anyon types) in our category. Because we have a concrete description of the simple objects in our category, it is possible to explicitly implement the braiding and fusion operations defined abstractly in [Oga22], and calculate those explicitly. We then show that the category we constructed is indeed equivalent to the one defined abstractly in [Oga22].
@@
The key difference between the present work and the abelian case studied in [Naa11, FN15] is the use of amplimorphisms, i.e. *-homomorphisms $\chi:\mathcal{A}\to M_{n}(\mathcal{A})$, instead of endomorphisms. [^1] For technical reasons we will in fact need to consider amplimorphisms of some slightly bigger algebra $\mathcal{B}$ containing $\mathcal{A}$. This can be understood as follows: recall that in the quantum double models, we can define ‘ribbon operators’ which create a pair of excitations from the ground state. To obtain single-anyon states, one sends one of the excitations off to infinity. For each irrep of ${\mathcal{D}}(G)$, there is a corresponding multiplet of ribbon operators, transforming according to the irrep, with the total number of operators in the multiplet given by the dimension of the irrep. Hence for non-abelian representations, one has more than one ribbon operator, which combine naturally into an amplimorphism.
@@
Although it is possible to pass from amplimorphisms to the endomorphisms used in [FN15, Oga22], as we shall see later, doing so requires making some choices, and one loses the explicit description of the map. Hence to identify the full superselection theory, we work mainly in the amplimorphism picture. In particular, we show that the amplimorphisms constructed in [Naa15] can be endowed with a tensor product and a braiding, analogous to the tensor product and braiding of endomorphisms in the DHR theory. More precisely, we construct a braided $\rm C^{*}$}-tensor category $\mathbf{Amp}$ of localized and transportable amplimorphisms, which includes as objects the amplimorphisms constructed in [Naa15]. We then consider the full subcategory $\mathbf{Amp}_{f}$ of $\mathbf{Amp}$ whose objects $\chi$ have finite dimensional $\mathrm{Hom}(\chi,\chi)$. This category can be shown to be semi-simple and closed under the monoidal product on $\mathbf{Amp}$, and we study the fusion rules (how tensor products decompose into irreducible objects) and the braiding. The result is that the category $\mathbf{Amp}_{f}$ of such amplimorphisms is equivalent to $\mathbf{Rep}_{f}{\mathcal{D}}(G)$ as braided tensor categories. Using the classification result of anyon sectors in this model obtained by two of the authors [BV25], it then follows that the list of constructed anyon sectors is a complete list of representatives of irreducible anyon sectors. This then completes the classification.
@@
A similar approach using amplimorphisms was taken in [SV93, NS97] to analyze topological defects of certain 1D quantum spin systems. In their setting the anyon sectors are localized in finite intervals, with the corresponding algebra of observables localized in that region being finite dimensional. This necessitated the use of amplimorphisms instead of endomorphisms. In our case localization is in infinite cone regions, and the situation is different. In particular, the unitary operators that can move the localization regions around no longer live in the quasi-local algebra $\mathcal{A}$. From a technical point of view this means that we cannot restrict to a purely $\rm{C}^{*}$-algebraic approach with operators in the quasi-local algebra (or suitable amplifications) only, but have to consider von Neumann algebras as well, in particular the cone algebras $\pi_{0}(\mathcal{A}_{\Lambda})^{\prime\prime}$. [^2] This is already true for the abelian case, it is not specific to the non-abelian model. These cone algebras are “big enough” in the sense that they are properly infinite [Naa12, FN15, Tom+24]. This allows us to directly relate the localized and transportable amplimorphisms to localized and transportable endomorphisms of some suitably defined auxiliary algebra, making the connection with the usual DHR theory in terms of endomorphisms.
@@
The paper is outlined as follows. In Section 2 we define the quantum double model and the associated categories of localized and tranportable amplimorphisms $\mathbf{Amp}$ and endomorphisms $\mathbf{DHR}$, as well as their ‘finite’ versions $\mathbf{Amp}_{f}$ and $\mathbf{DHR}_{f}$. We then state our main theorem, namely that the categories $\mathbf{Amp}_{f}$ and $\mathbf{DHR}_{f}$ are braided $\rm C^{*}$-tensor categories, equivalent to $\mathbf{Rep}_{f}{\mathcal{D}}(G)$. Section 3 is devoted to spelling out the braided $\rm C^{*}$-tensor structure of $\mathbf{Amp}$ and $\mathbf{DHR}$. These two categories are then shown to be equivalent in Section 4. Explicit localized and tranportable amplimorphisms corresonding to representations of ${\mathcal{D}}(G)$ are constructed in Section 5 by taking limits of ‘ribbon multiplets’. These explicit amplimorphisms are organized into full subcategories $\mathbf{Amp}_{\rho}$ of $\mathbf{Amp}$ for a fixed half-infinite ribbon $\rho$, which are later shown to be equivalent to $\mathbf{Amp}_{f}$. This section also establishes the key properties of these ribbon multiplets that underlie the fusion and braiding structure of $\mathbf{Amp}_{f}$. In Section 6 we rephrase the main result of [BV25], namely that the amplimorphisms corresponding to irreducible representations of ${\mathcal{D}}(G)$ constructed in the previous section exhaust all simple objects of $\mathbf{Amp}$. Together with semi-simplicity of $\mathbf{Amp}_{f}$, this implies that the $\mathbf{Amp}_{\rho}$ are full and faithful subcategories of $\mathbf{Amp}_{f}$. Finally, Section 7 proves the main theorem. The appendices collect well-known facts about ribbon operators and some technical results related to taking their limits.
@@
Copyright statement: For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising.
Data availability: We do not analyse or generate any datasets, because our work is entirely within a theoretical and mathematical approach.
Conflict of interests: The authors have no competing interests to declare that are relevant to the content of this article.
@@ 2.1 The quantum double model and its ground state
We first recall the definition of the quantum double model [Kit03] and introduce our notation. Throughout the paper, we fix a finite group $G$. Let $\Gamma$ be the triangular lattice in ${\mathbb{R}}^{2}$ and denote by $\Gamma^{E}$ the collection of oriented edges of $\Gamma$ which are oriented towards the right, see Figure 1.333We use the triangular lattice for simplicity, and to work in the same setting as [BV25], but believe the results hold for more general planar graphs as well. Denote by $\Gamma^{V},\Gamma^{F}$ the set of vertices and faces of $\Gamma$ respectively. To each edge $e\in\Gamma^{E}$ we associate a degree of freedom ${\mathcal{H}}_{e}\simeq{\mathbb{C}}[G]$ with basis $\{|g\rangle_{e}\,:\,g\in G\}$ and corresponding algebra ${\mathcal{A}}_{e}=\mathrm{End}({\mathcal{H}}_{e})\cong M_{|G|}(\mathbb{C})$. We define in the usual way local algebras of observables ${\mathcal{A}}^{\mathrm{loc}}_{X}$ supported on any $X\subset\Gamma^{E}$ and their norm closures ${\mathcal{A}}_{X}:=\overline{{\mathcal{A}}_{X}^{\mathrm{loc}}}^{\|\cdot\|}$. We write ${\mathcal{A}}={\mathcal{A}}_{\Gamma^{E}}$ and ${\mathcal{A}}^{\mathrm{loc}}={\mathcal{A}}_{\Gamma^{E}}^{\mathrm{loc}}$.
@@
The following theorem is proven in various sources [FN15, Cui+20, Tom+24, BV25].
@@ 2.2 Cone algebras, Haag duality, and the allowed algebra
We remark that all these cone algebras are properly infinite factors [Naa12, Oga22]. We will moreover assume that Haag duality holds for cones.
@@
Haag duality for cones is proven in [FN15] in the case $G$ is an abelian group. We believe the proof methods can be extended to the non-abelian case, however the analysis becomes considerably more technical since in the non-abelian case not all irreducible representations of the quantum double ${\mathcal{D}}(G)$ are one-dimensional anymore. In addition, a proof of Haag duality for a wide class of 2D quantum spin systems has been announced recently [OPR25], including in particular for the non-abelian quantum double models considered here.
Finally, we comment on the role that Haag duality plays. One can still construct the category of representations of superselection sectors, and show that the (equivalence classes of) irreducible representations are in one-to-one correspondence with the irreducible representations of ${\mathcal{D}}(G)$ [BV25]. By using this equivalence of categories the braided monoidal structure from $\mathbf{Rep}_{f}{\mathcal{D}}(G)$ can be transported to the category of superselection sectors. Haag duality is used to show that this in fact for example gives the natural braiding obtained from the Doplicher–Haag–Roberts approach. That is, it has the correct physical interpretation. Without Haag duality one can only do this for certain explicitly constructed representatives of each sector. [^4] This is the category $\mathbf{Amp}_{\rho}$ that we will define later. In this case, one can also explicitly construct the morphisms in the category as weak (or strong) operator limits of observables localized in some cone. This gives enough control over the localization of these intertwiners, which requires Haag duality in general. Using the explicit construction of the objects in the category, it can be directly checked that it is closed under the monoidal product of simple objects, and one can take finite direct sums. However, this analysis only works for the amplimorphisms constructed explicitly, and does not extend to arbitrary ampli (or endo-)morphisms, even if they are in the same superselection sector. For this reason, we prefer to assume (strict) Haag duality for cones to avoid making the analysis more technical than necessary.
@@
is the $\rm C^{*}$-algebra generated by the cone algebras of allowed cones. Note that the set of allowed cones is a directed set for the inclusion relation. Because we assume strict Haag duality, our algebra $\mathcal{B}$} is the same as what is denoted by $\mathfrak{B}_{(\theta,\phi)}$ in [Oga22, Eq. (2.5)] for suitable $(\theta,\phi)$. If only approximate Haag duality holds, it can be replaced with the definition there.
@@ 2.3 Categories of amplimorphisms and endomorphisms
We largely follow the notation and terminology of [SV93]. A *-homomorphism $\chi:{\mathcal{B}}\rightarrow M_{n\times n}({\mathcal{B}})$ is called an amplimorphism of degree nn. [^5] One can take amplifications with infinite dimensional Hilbert spaces, but for our purposes it is enough to consider only the case where $n$ is finite. We do not require such amplimorphisms to be unital. Given two amplimorphisms $\chi$ and $\chi^{\prime}$ of degrees $n$ and $n^{\prime}$ respectively, we let
@@
In the algebraic description of anyons, it is commonly assumed that all anyons have a conjugate (see for example  [Wan10, Sect. 6.3]), meaning that each anyon type can fuse to the vacuum with some conjugate type.
The assumption that an object in a $\rm C^{*}$-tensor category has a conjugate implies that it has a finite-dimensional endomorphism space [LR97, Lemma 3.2]. This is another way to see the necessity of restricting our attention to $\mathbf{Amp}_{f}$ if we want to show equivalence with $\mathbf{Rep}_{f}{\mathcal{D}}(G)$. Indeed, all finite dimensional representations of ${\mathcal{D}}(G)$ have conjugates.
@@
$\mathbf{DHR}$ is a braided $\rm C^{*}$-tensor subcategory of $\mathbf{Amp}$, see Section 3. We show in Section 7 that $\mathbf{DHR}_{f}$ is closed under the monoidal product of $\mathbf{DHR}$ and therefore inherits the braided $\rm C^{*}$-tensor structure. The category $\mathbf{DHR}$ is equivalent to the category $\mathcal{O}_{\Lambda_{0}}$ defined in [Oga22, Sect. 6]. One can think of $\mathcal{O}_{\Lambda_{0}}$ as the subcategory of $\mathbf{DHR}$ restricted to endomorphisms localized in a specific cone $\Lambda_{0}$, however by the transportability requirement, one sees that this is equivalent to $\mathbf{DHR}$ (compare with Sect. 7 here).